OpenStudy (anonymous):

Anyone else think that this is totally awesome?\[e^{i\pi}+1=0\]

myininaya (myininaya):

\[e^{ix}=\cos(x)+i \sin(x) => e^{i \pi}=\cos(\pi)+i \sin(\pi)=-1+i (0)=-1\] yes its okay lol

OpenStudy (jamesj):

Totally